MICROWAVE SPECTRUM OF THE FCO RADICAL IN THE ^{2}A′ ELECTRONIC GROUND STATE

\underline{HIDETA HABARA}^{a} and SATOSHI YAMAMOTO, Department of Physics and Research Center for the Early Universe, The University of Tokyo, Bunkyo-ku, Tokyo 113-0033, Japan.

The 1_{01}–0_{00}, 2_{02}–1_{01}, and 3_{03}–2_{02} rotational transitions of the FCO radical are observed at 22.3 GHz, 44.5 GHz, and 66.8 GHz, respectively, using a Fourier transform millimeter-wave spectrometer with a pulsed discharge nozzle. The FCO radical is produced by discharging F_{2}CO diluted in the Ar buffer gas. Twelve fine and hyperfine components for the three transitions are observed, and the effective rotational constant, the centrifugal distortion constant, the spin-rotation constant with its centrifugal correction term, and three hyperfine constants are determined. Furthermore the vibrational satellites for the ν_{1}, ν_{2}, ν_{3}, and 2ν_{3} states are also observed in the 22.3 GHz region. From the dipolar interaction constants, the principal axis of the dipolar interaction tensor is estimated, and is discussed in relation to the distribution of the unpaired electron.

aPresent adress, NASDA post doctral fellow, Institute for astrophysics and planetary sciences, Ibaraki University, 2-1-1 Bunkyo, Mito 310-8512, Japan